OpenStudy (shelby1290):

14-x(x+3)=2x-x(x-6)+8

Directrix (directrix):

What is the question - Solve for x? @shelby1290

OpenStudy (shelby1290):

@Directrix yes

Directrix (directrix):

Let's do this in parts: Multiply this out and simplify: 14-x*(x+3)= ? If you do that, then I'll share the load with you and I will crank out the right side. Okay?

OpenStudy (shelby1290):

okay

OpenStudy (shelby1290):

wait when you mutiply x by x is it 2x or x^2

Directrix (directrix):

x plus x = 2x x times x = x²

OpenStudy (shelby1290):

okay

OpenStudy (shelby1290):

when i did the first part i got: 14-x^2+3

OpenStudy (shelby1290):

oops 3x

Directrix (directrix):

Okay.

Directrix (directrix):

@shelby1290 Look at this again: 14-x*(x+3)= Should the 3x by +3x or -3x ?

Directrix (directrix):

Don't forget the Distributive Property ^^^ attachment review.

Directrix (directrix):

I'm starting on my part while you check yours.

OpenStudy (shelby1290):

wait what are you asking me to change ?

Directrix (directrix):

I'm saying that you have a bad sign on the 3x. when i did the first part i got: 14-x^2+3x You need to correct it or we will get a wrong answer in the end.

Directrix (directrix):

You messed up when you multiplied this: -x*(x+3) The -x is to be distributed over the x AND the + 3.

Directrix (directrix):

2x-x(x-6)+8 = 2x - x*x -x*(-6) + 8 = 2x - x² + 6x + 8 = -x² + 8x + 8 --> This is the right side.

Directrix (directrix):

For the left side, you were supposed to get: -x² - 3x + 14

OpenStudy (shelby1290):

wait but how come you moved 14

Directrix (directrix):

14-x(x+3)=2x-x(x-6)+8 -x² - 3x + 14 = -x² + 8x + 8 I put my work on the right and yours on the left above. The task now is to solve this equation. -x² - 3x + 14 = -x² + 8x + 8 +x² +x² ----------------------------- @shelby1290 What do we get now after adding x² to both sides.

OpenStudy (shelby1290):

dont you collect like terms or is this something else

Directrix (directrix):

>>so the left side will be 14-x^2-3x ? Yes. >>wait but how come you moved 14 ? It was not required. I put the terms in descending order of the powers. You can leave the 14 as you had it. But, the mistake with the 3 and the x had to be corrected.

Directrix (directrix):

>>dont you collect like terms Right.

OpenStudy (shelby1290):

i believe -x^2 and -x^2 would cancel out ?? ( not sure) 3x+8x=11x or 3x-8x =-5x sorry but i just suck at this unit and math in general :/

Directrix (directrix):

-x² - 3x + 14 = -x² + 8x + 8 +x² +x² --------------------------- -3x + 14 = 8x + 8 -8x -8x ------------------------ -11x + 14 = 8 - 14 -14 ------------------------ -11x = -6 x = ? @shelby1290

Directrix (directrix):

@shelby1290 What is the value of x?

OpenStudy (shelby1290):

x=5

Directrix (directrix):

Please post the work for that x = 5. Okay?

OpenStudy (shelby1290):

to get 5 i did -6+11

OpenStudy (shelby1290):

am i right ?

Directrix (directrix):

No. The -11 is not really -11 but is -11x. -11x = -6 Divide both sides by -11 x = (-6)/(-11) = 6/11

OpenStudy (shelby1290):

0.54

OpenStudy (shelby1290):

do you simplify it to a decimal or leave it as a fraction

Directrix (directrix):

I would leave the answer as 6/11 but I don't know what you have been instructed to do regarding decimals and approximation.

Directrix (directrix):

If you have answer options, then look at them. If not, then I don't know but would leave the answer at x = 6/11.

OpenStudy (shelby1290):

okay thank you so much @Directrix

Directrix (directrix):

You are welcome.